# LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues

### v1

@@ Abstract.
Multi-turn dialogues are essential in many real-world applications of large language models, such as chatbots and virtual assistants. As conversation histories become longer, existing large language models face increasing computational and memory challenges, which hinder their ability to provide efficient and responsive interactions. Most current acceleration methods either compress the context or optimize key value caching, but they often rely on fixed or position-based heuristics that do not adapt well to the dynamic and unpredictable patterns found in actual multi-turn conversations. In this paper, we present LoopServe, an adaptive dual-phase inference acceleration framework for large language models in multi-turn dialogues. LoopServe introduces two main innovations. First, it performs online sparsification during the prefilling phase by dynamically selecting the most important parts of the attention matrix for each new input. Second, it uses progressive key value compression during decoding by adaptively maintaining a relevant and efficient cache based on the most recently generated output tokens. We also propose a new benchmark with eleven multi-turn datasets that reflect realistic query positions and conversational dependencies. Extensive experiments demonstrate that LoopServe consistently achieves superior effectiveness compared to existing baselines and significantly accelerates LLM inference across a wide range of long-context dialogue tasks.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025. doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097. doi:XX.XX/XXX.XX
1. Introduction
@@
Multi-turn dialogues are at the core of numerous real-world applications, from customer service chatbots to virtual assistants and collaborative agents. These scenarios demand that large language models (LLMs) (Hadi et al., 2023; Deng et al., 2025; Li et al., 2024c; Zhou et al., 2024; van Renen et al., 2024) not only generate coherent responses but also maintain contextual consistency across lengthy, evolving conversations. As the number of dialogue turns increases, so does the computational workload. For instance, processing a multi-turn dialogue comprising 10,000 tokens with Llama-3.1-70B (Grattafiori et al., 2024) can demand trillions of floating-point operations (FLOPs), quickly challenging the limits of real-time inference. Despite the remarkable progress of LLMs such as GPT (Brown et al., 2020; Radford et al., 2018, 2019), Llama (Grattafiori et al., 2024; Touvron et al., 2023), and DeepSeek (DeepSeek-AI, 2024, 2025), their inefficiency in handling multi-turn dialogues remains largely unaddressed.
@@
In a typical multi-turn setting, each new user turn expands the conversation history, requiring the LLM to process ever-growing input sequences. The model’s self-attention mechanism, which lies at the heart of the Transformer (Vaswani, 2017), needs to compute pairwise attention scores between every pair of tokens in the input. Specifically, given an input sequence of length $n$, an LLM with $P$ parameters and a hidden dimension $d$, the time complexity of generating $m$ tokens is $\mathcal{O}(m((n+m)^{2}d+P))$. For instance, in a 3-turn conversation with 5000 tokens per turn, the effective context length for the model reaches 15,000 tokens, resulting in quadratic growth in both computational cost and memory usage. This compounding effect of context accumulation makes real-time, cost-efficient inference increasingly difficult as the conversation progresses.
This challenge is distinct from single-turn tasks, where the input context is static and the user query is typically placed at the end of the sequence. In contrast, multi-turn dialogues feature dynamic context accumulation and variable query positions: user questions may appear at the beginning, middle, or end of the input. This variability causes the attention patterns within the model to change unpredictably across turns and inputs, requiring the model to flexibly maintain and retrieve relevant information from any point in the conversation. The resulting attention matrices not only grow in size with each turn but also exhibit highly dynamic and input-dependent sparsity patterns, further exacerbating the inefficiency of existing inference methods.
@@
Recent research on LLM inference acceleration has primarily focused on two directions: context compression and Key-Value (KV) compression. Firstly, Context compression techniques, such as LLMLingua (Pan et al., 2024), CompAct (Yoon et al., 2024), and RAG-based approaches (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024; Wang et al., 2025), reduce input length by filtering, summarizing, or structuring the context. While effective in lowering computational costs, these strategies often sacrifice fine-grained details, which can degrade output fidelity. Secondly, KV compression approaches aim to reduce the computational burden of attention weight calculations during both the prefilling and decoding stages. In the prefilling stage, where the attention matrix is computed for all token pairs, methods (Jiang et al., 2024a; Lv et al., 2024; Lai et al., 2025), such as Minference (Jiang et al., 2024a), sparsify the attention matrix to reduce quadratic computation (See Table 2). In the decoding stage, KV caches store precomputed Key and Value vectors for reuse, minimizing redundant computations during autoregressive text generation. For example, H2O (Zhang et al., 2023b), PQCache (Zhang et al., 2025), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) select or cache only the most relevant tokens based on attention weights or fixed token positions. However, these methods predominantly rely on static or position-based heuristics, which are ill-suited for the dynamic, input-dependent patterns observed in real-world multi-turn dialogues.
@@
Moreover, current evaluation benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023) for LLM acceleration misrepresent real-world dialogue scenarios. Most benchmarks (Bai et al., 2024a; Li et al., 2025; Dacheng Li* and Zhang, 2023) assume queries are always placed at the end of the input and focus on single-turn tasks, which oversimplifies the problem and favors acceleration methods that exploit positional biases. As a result, approaches (Zhang et al., 2023b; Li et al., 2024b; Feng et al., 2024) that perform well on these benchmarks often fail to generalize to realistic dialogue scenarios where queries may appear at arbitrary positions and contextual dependencies span multiple turns.
To address these challenges, we conduct experiments and summarize three key findings: (1) The sparsity patterns of attention matrices are highly dynamic across inputs, heads, and layers. A small fraction of vertical and slash lines in the attention matrix accounts for most attention weights, but these patterns vary significantly and cannot be effectively captured by static or offline sparsification methods. (2) The position of questions in user queries significantly affects the performance of existing LLM acceleration techniques, which assume queries are always at the end of the input sequence. However, queries often appear in the middle or beginning of the context. (3) Token importance evolves dynamically during multi-turn dialogues as new outputs are generated. Fixed token selection strategies fail to capture these changing dependencies, leading to inefficiencies and reduced conversational quality.
@@
Based on these observations, we propose LoopServe, an adaptive dual-phase LLM inference acceleration framework specifically designed for multi-turn dialogues. LoopServe introduces two key innovations: online prefilling sparsification and progressive KV compression. In the prefilling phase, LoopServe dynamically identifies and selects the most critical components of the attention matrix, specifically the vertical and slash lines that account for the majority of attention weights. Unlike fixed sparsification methods, LoopServe adapts online, ensuring computational efficiency while maintaining high attention fidelity. In the decoding phase, LoopServe employs a progressive KV compression strategy. It dynamically selects and compresses relevant input tokens based on the most recently generated output tokens. This approach ensures the KV cache remains efficient and relevant throughout the decoding process, significantly reducing computational overhead without compromising output quality. Also, we propose a multi-turn long-context benchmark, including 11 datasets. This benchmark captures diverse query positions and multi-turn dependencies, providing a more realistic evaluation framework for multi-turn dialogue scenarios. The contributions of this paper are summarized as follows:
@@
  • We empirically demonstrate that attention patterns and key token positions in multi-turn dialogues are highly dynamic and unpredictable, revealing the limitations of static sparsification and KV selection methods.
@@
  • We introduce LoopServe, a novel dual-phase LLM inference acceleration system that integrates online attention sparsification and progressive KV compression, substantially improving efficiency and robustness for multi-turn dialogue inference.

### v2

@@ Abstract
Multi-turn dialogues are essential in many real-world applications of large language models, such as chatbots and virtual assistants. As conversation histories become longer, existing large language models face increasing computational and memory challenges, which hinder their ability to provide efficient and responsive interactions. Most current acceleration methods either compress the context or optimize key value caching, but they often rely on fixed or position-based heuristics that do not adapt well to the dynamic and unpredictable patterns found in actual multi-turn conversations. As a result, these models cannot accurately identify and prioritize the most relevant context, leading to degraded response quality. In this paper, we present LoopServe, an adaptive dual-phase inference acceleration framework for large language models in multi-turn dialogues. LoopServe introduces two main innovations. First, it performs online sparsification during the prefilling phase by dynamically selecting the most important parts of the attention matrix for each new input. Second, it uses progressive key value compression during decoding by adaptively maintaining a relevant and efficient cache based on the most recently generated output tokens. We also propose a new benchmark with eleven multi-turn datasets that reflect realistic query positions and conversational dependencies. Extensive experiments demonstrate that LoopServe consistently achieves superior effectiveness compared to existing baselines and significantly accelerates LLM inference across a wide range of long-context dialogue tasks.
1 Introduction
@@
Multi-turn dialogues are at the core of numerous real-world applications, from customer service chatbots to virtual assistants and collaborative agents. These scenarios demand that large language models (LLMs) (Hadi et al., 2023; Deng et al., 2025; Li et al., 2024a; Zhou et al., 2024; van Renen et al., 2024) not only generate coherent responses but also maintain contextual consistency across lengthy, evolving conversations. As the number of dialogue turns increases, so does the computational workload. For instance, processing a multi-turn dialogue comprising 10,000 tokens with Llama-3.1-70B (Grattafiori et al., 2024) can demand trillions of floating-point operations (FLOPs), quickly challenging the limits of real-time inference. Despite the remarkable progress of LLMs such as GPT (Brown et al., 2020; Radford et al., 2018; 2019), Llama (Grattafiori et al., 2024; Touvron et al., 2023), and DeepSeek (DeepSeek-AI, 2024; 2025), their inefficiency in handling multi-turn dialogues remains largely unaddressed.
@@
In a typical multi-turn setting, each new user turn expands the conversation history, requiring the LLM to process ever-growing input sequences. The model’s self-attention mechanism, which lies at the heart of the Transformer (Vaswani, 2017), needs to compute pairwise attention scores between every pair of tokens in the input. Specifically, given an input sequence of length $n$, an LLM with $P$ parameters and a hidden dimension $d$, the time complexity of generating $m$ tokens is $\mathcal{O}(m((n+m)^{2}d+P))$. For instance, in a 3-turn conversation with 5000 tokens per turn, the effective context length for the model reaches 15,000 tokens, resulting in quadratic growth in both computational cost and memory usage. This compounding effect of context accumulation makes real-time, cost-efficient inference increasingly difficult as the conversation progresses. Different from single-turn tasks, the context in multi-turn dialogues accumulates dynamically and queries may appear at the beginning, middle, or end of the input, causing attention patterns to shift unpredictably. The resulting attention matrices not only grow with each turn but also exhibit highly dynamic and input-dependent sparsity, exacerbating the inefficiency of current inference methods.
@@
Recent research proposes accelerating LLM inference by reducing the computational burden of attention weight calculations during both the prefilling and decoding stages. In the prefilling stage, where the attention matrix is computed for all token pairs, methods (Jiang et al., 2024a; Lv et al., 2024; Lai et al., 2025), such as Minference (Jiang et al., 2024a), use fixed pattern to sparsify the attention matrix to reduce quadratic computation. During decoding, KV caches store precomputed Key and Value vectors to reduce redundant computation. Methods like H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) cache tokens selected based on tokens at the end of the query. However, these approaches rely on static or position-based heuristics and cannot adapt to the dynamic, input-dependent patterns of real multi-turn dialogues.
@@
Moreover, current evaluation benchmarks (Li et al., 2024a; 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023) for LLM acceleration misrepresent real-world dialogue scenarios. Most benchmarks (Bai et al., 2024a; Li et al., 2025; Dacheng Li* & Zhang, 2023) assume queries are always placed at the end of the input and focus on single-turn tasks, which oversimplifies the problem and favors acceleration methods that exploit positional biases. As a result, approaches (Zhang et al., 2023; Li et al., 2024b; Feng et al., 2024) that perform well on these benchmarks often fail to generalize to realistic dialogue scenarios where queries may appear at arbitrary positions and contextual dependencies span multiple turns.
@@
In this paper, we propose LoopServe, an adaptive dual-phase LLM inference acceleration framework specifically designed for multi-turn dialogues. LoopServe features two core innovations: online prefilling sparsification and progressive KV compression. In the prefilling phase, LoopServe dynamically identifies and selects the most critical components of the attention matrix, focusing on the vertical and slash line patterns that contribute most to attention weights. Unlike fixed sparsification methods, LoopServe adapts in real time to maintain both efficiency and high attention fidelity. During decoding, LoopServe applies progressive KV compression by dynamically selecting and compressing relevant input tokens based on the most recently generated outputs. This strategy keeps the KV cache efficient and relevant throughout decoding, significantly reducing computational overhead without compromising output quality. We also introduce a multi-turn long-context benchmark containing 11 datasets. This benchmark captures diverse query positions and multi-turn dependencies, offering a more realistic evaluation framework for dialogue scenarios. The contributions of this paper are summarized as follows:
@@
  • We empirically reveal that attention patterns and key token positions in multi-turn dialogues are highly dynamic, limiting static sparsification and KV selection.
@@
  • We present LoopServe, a dual-phase LLM acceleration framework with online attention sparsification and progressive KV compression, improving multi-turn inference efficiency.